OpenStudy (anonymous):

Walt and Amy each have the same number of pennies. Walt made equal stacks of 10 pennies with none left over. Amy made equal stacks of 8 pennies with none left over. What is the least number of pennies that each person could have? Explain how you found your answer.

OpenStudy (anonymous):

intuitively - 50

OpenStudy (gsx):

40...

OpenStudy (anonymous):

50 = 6*8 + 2

OpenStudy (gsx):

There should be no leftovers... 5*8=40; 4*10=40

OpenStudy (anonymous):

if it was 40 that would be 5 stacks with none left over

OpenStudy (anonymous):

oh sorry - my eys are playing tricks - i took it to be 2 left over!!

OpenStudy (anonymous):

yup - 40 is correct

OpenStudy (anonymous):

ok thank you!!! :)

OpenStudy (anonymous):

yw - i'd better have an eye test lol!!

OpenStudy (anonymous):

lol